A game is set up so that if a face card is drawn from a fair deck, $50 is won. If a card value less than 6 is drawn, $10 is lost. If a 9 or a ten is drawn, $0.50 is won. Any other card drawn yields no gain nor loss. How much should the house charge to draw a card if the expected value of the game needs to be zero (the house breaks even)?
Answer #1

Ans:

Probability of drawing a face card=12/52

Probability of drawing a card value less than 6=20/52

Probability of drawing a 9 or 10=8/52

For break even,

C=(12/52)*50-(20/52)*10+(8/52)*0.50

C=7.77 dollars